Circular No. 7228

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999df IN CGCG 274-026 T. Puckett and I. Siegert, Mountain Town, GA, report the discovery

of an apparent supernova (mag 18.2) on an unfiltered CCD frame (limiting mag 20.0) taken with a 0.60-m automated supernova patrol telescope on July 20.16 UT. SN 1999df is located at α = 15^{h}08^{m}56^{s}.52, δ = +52^{o}26′53′′.1 (equinox 2000.0), which is 5′′.9 north of CGCG 274-026. The new object was also present on an unfiltered CCD frame taken on July 21.09, but it was not present on an unfiltered CCD frame taken on June 8.17 (limiting mag ∼ 20.5), nor does it appear on a second-generation Palomar Sky survey image (limiting mag 20.7) taken on 1991 April 19. An image from Puckett

can be accessed at http://www.cometwatch.com/sn pgc 54081.jpg.

NO SUPERNOVA 1999db T. Matheson, M. Modjaz, and A. V. Filippenko, University of Califor-

nia at Berkeley, write: “A CCD spectrum (range 420–700 nm) obtained on July 18 with the Shane 3-m reflector at Lick Observatory shows that SN 1999db (cf. IAUC 7227) is probably a quasi-stellar object with redshift 1.9. Prominent emission at 450 nm is identified with C IV 155-nm, and a much weaker line near 550 nm could be C III] 191-nm. In any case, the spectrum does not resemble those of known supernovae.”

V1493 AQUILAE = NOVA AQUILAE 1999 N. James, Chelmsford, England, reports a precise CCD position for

this nova using a 0.3-m reflector on July 14.919 UT: α = 19^{h}07^{m}36^{s}.91, δ = +12^{o}31′26′′.7 (equinox 2000.0; eleven GSC stars, rms residuals 0′′.22 in α, 0′′.23 in δ). G. Masi, Ceccano, Italy, reports the following position end figures for the nova, measured from six USNO A2.0 reference stars: 36^{s}.89, 26′′.5 (mean residuals 0′′.04 in α, 0′′.10 in δ).

CCD magnitudes (Czech data based on comparison star GSC 1048.863, with assumed V = 8.40 ± 0.02, B − V = +0.31 ± 0.02 from Tycho Catalogue): July 14.955 UT, V = 10.16 ± 0.01 (M. Zejda, Brno, Czech Rep.); 15.820, V = 10.59 ± 0.01, B − V = +0.40 ± 0.02 (J. Šafář, Brno, Czech Rep.); 16.911, R = 8.55 (G. Masi, Ceccano, Italy); July 15.967, V = 10.64 ± 0.02 (P. Hájek, Vyškov, Czech Republic); 17.020, V = 11.10 ± 0.02 (Hájek); 17.035, V = 11.02 ± 0.03 (D. Hanžl, Brno, Czech Rep.); 17.999, V = 11.27 ± 0.02, B − V = +0.24 ± 0.03 (Šafář); 18.000, V = 11.20 ± 0.03 (Hájek); 18.890, V = 11.42 ± 0.01, B − V = +0.39 ± 0.04 (Zejda).

1999 July 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green